# 0-supple.tex

### v1

@@ C.2 Proof of the Results for the Heterogeneous Setting
$\sup_{j,\ell}\left\lVert\widehat{V}_{j,\ell}\widehat{V}_{j,\ell}^{\top}-V_{j,\ell}V_{j,\ell}^{\top}\right\rVert_{\rm F}$ $=O_{\mathbb{P}}\left(\frac{\sqrt{pr_{j,V,\ell}}\lambda_{\max}\sigma+\sqrt{prr_{j,V,\ell}/L}\lambda_{\max}^{2}\kappa\sigma/\Delta+\frac{\sqrt{r_{j,V,\ell}}pr\lambda_{\max}^{2}(\kappa^{2}+\kappa_{0}^{2})\sigma^{2}}{\Delta^{2}}}{\lambda_{\min}^{2}}\right)$
$=O_{\mathbb{P}}\left(\frac{\sqrt{pr_{V}}\kappa_{0}\sigma(1+\kappa^{2}\sqrt{r/L})}{\lambda_{\min}}+\frac{\sqrt{r_{V}}pr\kappa^{2}(\kappa^{2}+\kappa_{0}^{2})\sigma^{2}}{\lambda_{\min}^{2}}\right).$
Proof of Theorem 4.1.
Define $\Sigma_{j}^{*}=\sum_{\ell=s,t}w_{\ell}\mathbb{E}\left[\widehat{U}_{j,\ell}\widehat{U}_{j,\ell}^{\top}\right]$ and let $U_{j}^{*}=\mathrm{svd}_{r_{j,U}}(\Sigma_{j}^{*})$. By the proof of Theorem 3.1, we have that
$\sup_{j}\left\lVert\widehat{U}_{j}\widehat{U}_{j}^{\top}-U_{j}^{*}U_{j}^{*\top}\right\rVert_{\rm F}$ $\lesssim\quad\left\lVert\sum_{\ell=s,t}w_{\ell}(1-\mathbb{E})\left[\widehat{U}_{j,\ell}\widehat{U}_{j,\ell}^{\top}\right]\right\rVert_{\rm F}$
$=O_{\mathbb{P}}\left(\sqrt{\sum_{\ell=s,t}w_{\ell}^{2}\left\lVert(1-\mathbb{E})\left[\widehat{U}_{j,\ell}\widehat{U}_{j,\ell}^{\top}\right]\right\rVert_{\rm F}^{2}}\right)$
$=O_{\mathbb{P}}\left(\frac{\sqrt{pr_{U}}\kappa\sqrt{w_{s}^{2}\sigma_{s}^{2}+w_{t}^{2}\sigma_{t}^{2}}}{\Delta}\right),$
@@
$\sup_{j}\left\lVert U_{j}^{*}U_{j}^{*\top}-U_{j}U_{j}^{\top}\right\rVert_{\rm F}\lesssim\frac{pr(\kappa^{2}+\kappa_{0}^{2})(w_{s}\sigma_{s}^{2}+w_{t}\sigma_{t}^{2})}{\Delta^{2}}.$
Therefore,
@@
$\sup_{j}\left\lVert\widehat{U}_{j}\widehat{U}_{j}^{\top}-U_{j}U_{j}^{\top}\right\rVert_{\rm F}=O_{\mathbb{P}}\left(\frac{\sqrt{pr}\kappa\sqrt{w_{s}^{2}\sigma_{s}^{2}+w_{t}^{2}\sigma_{t}^{2}}}{\Delta}+\frac{pr(\kappa^{2}+\kappa_{0}^{2})(w_{s}\sigma_{s}^{2}+w_{t}\sigma_{t}^{2})}{\Delta^{2}}\right).$
Using the same argument in the proof of Theorem 3.1, we have
$\quad\sup_{j}\left\lVert\widehat{V}_{j,t}\widehat{V}_{j,t}^{\top}-V_{j,t}V_{j,t}^{\top}\right\rVert_{\rm F}$
@@
$=O_{\mathbb{P}}\left(\frac{\sqrt{pr_{V}}\kappa_{0}}{\lambda_{\min}}\Big{(}\sigma_{t}+\kappa^{2}\sqrt{r}\sqrt{w_{s}^{2}\sigma^{2}_{s}+w_{t}^{2}\sigma_{t}^{2}}\Big{)}+\frac{\sqrt{r_{V}}pr\kappa^{2}(\kappa^{2}+\kappa_{0}^{2})(w_{s}\sigma_{s}^{2}+w_{t}\sigma_{t}^{2})}{\lambda_{\min}^{2}}\right).$
Since $\kappa_{0},\kappa=O(1)$, $\Delta/\max(\sigma_{s},\sigma_{t})\gtrsim\sqrt{pr}$ and
$(w_{s}\sigma_{s}^{2}+w_{t}\sigma_{t}^{2})\leq\sqrt{2}\max(\sigma_{s},\sigma_{t})\sqrt{w_{s}^{2}\sigma_{s}^{2}+w_{t}^{2}\sigma_{t}^{2}},$
@@
$\frac{pr(\kappa^{2}+\kappa_{0}^{2})(w_{s}\sigma_{s}^{2}+w_{t}\sigma_{t}^{2})}{\Delta^{2}}\lesssim\frac{\sqrt{pr}\kappa\sqrt{w_{s}^{2}\sigma_{s}^{2}+w_{t}^{2}\sigma_{t}^{2}}}{\Delta},$
@@
$\sup_{j}\left\lVert\widehat{U}_{j}\widehat{U}_{j}^{\top}-U_{j}U_{j}^{\top}\right\rVert_{\rm F}=O_{\mathbb{P}}\left(\frac{\sqrt{pr}\kappa\sqrt{w_{s}^{2}\sigma_{s}^{2}+w_{t}^{2}\sigma_{t}^{2}}}{\Delta}\right).$
Similarly, it holds that
$\frac{\sqrt{r_{V}}pr\kappa^{2}(\kappa^{2}+\kappa_{0}^{2})(w_{s}\sigma_{s}^{2}+w_{t}\sigma_{t}^{2})}{\lambda_{\min}^{2}}\lesssim\frac{\sqrt{pr_{V}}\kappa_{0}}{\lambda_{\min}}\Big{(}\kappa^{2}\sqrt{r}\sqrt{w_{s}^{2}\sigma^{2}_{s}+w_{t}^{2}\sigma_{t}^{2}}\Big{)},$
which yields that
$\sup_{j}\left\lVert\widehat{V}_{j,t}\widehat{V}_{j,t}^{\top}-V_{j,t}V_{j,t}^{\top}\right\rVert_{\rm F}=O_{\mathbb{P}}\left(\frac{\sqrt{pr_{V}}\kappa_{0}}{\lambda_{\min}}\Big{(}\sigma_{t}+\kappa^{2}\sqrt{r}\sqrt{w_{s}^{2}\sigma^{2}_{s}+w_{t}^{2}\sigma_{t}^{2}}\Big{)}\right).$
Plugging in $w_{s}=\frac{\sigma_{t}^{2}}{\sigma_{s}^{2}+\sigma_{t}^{2}}$ and $w_{t}=\frac{\sigma_{s}^{2}}{\sigma_{s}^{2}+\sigma_{t}^{2}}$ leads to
$\sup_{j}\left\lVert\widehat{U}_{j}\widehat{U}_{j}^{\top}-U_{j}U_{j}^{\top}\right\rVert_{\rm F}=O_{\mathbb{P}}\left(\frac{\sqrt{pr}\kappa\sigma_{s}\sigma_{t}}{\Delta\sqrt{\sigma_{s}^{2}+\sigma_{t}^{2}}}\right),$

### v3

@@ D.2 Proof of the Results for the Heterogeneous Setting
$\sup_{j,\ell}\left\lVert\widehat{V}_{j,\ell}\widehat{V}_{j,\ell}^{\top}-V_{j,\ell}V_{j,\ell}^{\top}\right\rVert_{\rm F}$ $\lesssim\frac{\sqrt{pr_{V}}\lambda_{\max}\sigma+\sqrt{prr_{V}/L}\lambda_{\max}^{2}\kappa_{0}\sigma/\Delta+\frac{\sqrt{r_{V}}pr\lambda_{\max}^{2}\kappa_{0}^{2}\sigma^{2}}{\Delta^{2}}}{\lambda_{\min}^{2}}$
$\lesssim\frac{\sqrt{pr_{V}}\kappa_{0}\sigma(1+\kappa\kappa_{0}\sqrt{r/L})}{\lambda_{\min}}+\frac{\sqrt{r_{V}}pr\kappa^{2}\kappa_{0}^{2}\sigma^{2}}{\lambda_{\min}^{2}},$
with probability at least $1-\widetilde{C}_{1}e^{-\widetilde{c}_{1}p}$. ∎
Proof for Theorem 3.2.
By the proof of Theorem 2.1, with probability at least $1-C_{1}^{\prime}Le^{-c_{1}^{\prime}p}\geq 1-\widetilde{C}_{1}e^{-\widetilde{c}_{1}p}$, we have that
$\left\lVert\overline{W}_{j}-U_{j}U_{j}^{\top}-\frac{1}{L}\sum_{\ell=1}^{L}V_{j,\ell}V_{j,\ell}^{\top}\right\rVert_{\rm F}\leq 2C_{2}^{\prime}\lambda_{\min}^{-1}\sigma\sqrt{pr}.$ (56)
[…]
$\lambda_{\max}(\overline{V}_{j})=\lambda_{\max}(\Sigma_{j,V})\leq 1-\frac{\eta_{V}}{3}.$
First, rewrite $\overline{V}_{j}=\frac{1}{L}\widetilde{V}_{j}\widetilde{V}_{j}^{\top}$, where $\widetilde{V}_{j}:=[V_{j,1}\,V_{j,2}\dots V_{j,L}]\in\mathbb{R}^{p_{j}\times\widetilde{r}_{j,V}}$ and $\widetilde{r}_{j,V}:=\sum_{\ell=1}^{L}r_{j,V,\ell}$. Note that
$I_{\widetilde{r}_{j,V}}-\frac{1}{L}\widetilde{V}_{j}^{\top}\widetilde{V}_{j}=\begin{pmatrix}\frac{L-1}{L}I_{r_{j,V,1}}&-\frac{1}{L}V_{j,1}^{\top}V_{j,2}&\cdots&-\frac{1}{L}V_{j,1}^{\top}V_{j,L}\\ -\frac{1}{L}V_{j,2}^{\top}V_{j,1}&\frac{L-1}{L}I_{r_{j,V,2}}&\cdots&-\frac{1}{L}V_{j,2}^{\top}V_{j,L}\\ &\cdots&\cdots&\\ -\frac{1}{L}V_{j,L}^{\top}V_{j,1}&-\frac{1}{L}V_{j,L}^{\top}V_{j,2}&\cdots&\frac{L-1}{L}I_{r_{j,V,L}}\\ \end{pmatrix}.$
Assume that $\lambda$ is an eigenvalue of $I_{\widetilde{r}_{j,V}}-\frac{1}{L}\widetilde{V}_{j}^{\top}\widetilde{V}_{j}$, and corresponding eigenvector is $\bm{v}$. Write $\bm{v}=(\bm{v}_{1}^{\top},\bm{v}_{2}^{\top},\dots,\bm{v}_{L}^{\top})^{\top}$, where $\bm{v}_{\ell}\in\mathbb{R}^{r_{j,V,\ell}}$. Let ~$\widetilde{\ell}=\operatorname*{\arg\!\max}_{\ell}\left\lVert\bm{v}_{\ell}\right\rVert_{2}$. Since $(I_{\widetilde{r}_{j,V}}-\frac{1}{L}\widetilde{V}_{j}^{\top}\widetilde{V}_{j})\bm{v}=\lambda\bm{v}$, we have
$\left\lVert\left(\lambda-\frac{L-1}{L}\right)\bm{v}_{\widetilde{\ell}}\right\rVert_{2}=\frac{1}{L}\left\lVert\sum_{\ell^{\prime}\neq\widetilde{\ell}}V_{j,\widetilde{\ell}}^{\top}V_{j,\ell^{\prime}}\bm{v}_{\ell^{\prime}}\right\rVert_{2}\leq\frac{1}{L}\sum_{\ell^{\prime}\neq\widetilde{\ell}}\left\lVert V_{j,\widetilde{\ell}}^{\top}V_{j,\ell^{\prime}}\right\rVert_{2}\left\lVert\bm{v}_{\ell^{\prime}}\right\rVert_{2}\leq\frac{(L-1)(1-\eta_{V})}{L}\left\lVert\bm{v}_{\widetilde{\ell}}\right\rVert_{2},$
where the last inequality follows from the definition that $\eta_{V}=\min_{j,\ell}\sum_{\ell^{\prime}\neq\ell}(1-\left\lVert V_{j,\ell}^{\top}V_{j,\ell^{\prime}}\right\rVert_{2})/(L-1)$ and thus ~$\sum_{\ell^{\prime}\neq\widetilde{\ell}}\left\lVert V_{j,\widetilde{\ell}}^{\top}V_{j,\ell^{\prime}}\right\rVert_{2}\leq(L-1)(1-\eta_{V})$. Therefore, we obtain that $\lambda\geq\frac{\eta_{V}(L-1)}{L}$, and this is true for all of the eigenvalues of $I_{\widetilde{r}_{j,V}}-\frac{1}{L}\widetilde{V}_{j}^{\top}\widetilde{V}_{j}$.
By Woodbury matrix identity, it holds that
$(I_{p_{j}}-\overline{V}_{j})^{-1}=I_{p_{j}}+\frac{1}{L}\widetilde{V}_{j}\left(I_{\widetilde{r}_{j,V}}-\frac{1}{L}\widetilde{V}_{j}^{\top}\widetilde{V}_{j}\right)^{-1}\widetilde{V}_{j}^{\top}.$
Since $\left\lVert\widetilde{V}_{j}\right\rVert_{2}\leq\sqrt{\sum_{\ell=1}^{L}\left\lVert V_{j,\ell}\right\rVert_{2}^{2}}\leq\sqrt{L}$, we obtain
$\left\lVert(I_{p_{j}}-\overline{V}_{j})^{-1}\right\rVert_{2}\leq 1+\frac{L}{\eta_{V}(L-1)},$
yielding that
$\left\lVert\overline{V}_{j}\right\rVert_{2}=1-\lambda_{\min}(I_{p_{j}}-\overline{V}_{j})=1-\left\lVert(I_{p_{j}}-\overline{V}_{j})^{-1}\right\rVert_{2}^{-1}\leq 1-\frac{\eta_{V}}{3}.$
[…]
$\sigma_{k}(\overline{W})\geq 1-2C_{2}^{\prime}\frac{\sigma\sqrt{pr}}{\lambda_{\min}},\forall k\leq r_{j,U},$
@@
$\sigma_{k}(\overline{W})\leq 1-\frac{\eta_{V}}{3}+2C_{2}^{\prime}\frac{\sigma\sqrt{pr}}{\lambda_{\min}},\forall k\geq r_{j,U}+1.$
Therefore, since $\frac{\sigma\sqrt{pr}}{\lambda_{\min}}=o(1)$and $\eta_{V}\geq\eta_{0}$, for any $\delta_{0}\in(0,\eta_{0}/3)$, we have $\mathbb{P}(\widehat{r}_{j,U}=r_{j,U})\geq 1-\widetilde{C}_{1}e^{-\widetilde{c}_{1}p}$.
D.3 Proof of the Results for the Knowledge Transfer Setting
Proof of Theorems 4.1 and C.1.
It suffices to show Theorem C.1 since Theorem 4.1 is a special case with $L=2$. Analogous to (51), (52), and (53), we have that
@@ D.3 Proof of the Results for the Knowledge Transfer Setting
$\sup_{j,\ell}\frac{1}{\sigma_{\ell}^{2}}\left\lVert\widehat{U}_{j,\ell}\widehat{U}_{j,\ell}^{\top}-U_{j}U_{j}^{\top}-f_{j,\ell}\big{[}\zeta_{j,\ell,1}U_{j}\big{]}U_{j}^{\top}-f_{j,\ell}\big{[}\zeta_{j,\ell,1}^{\top}U_{j}\big{]}U_{j}^{\top}-U_{j}f_{j,\ell}\big{[}\zeta_{j,\ell,1}U_{j}\big{]}^{\top}-U_{j}f_{j,\ell}\big{[}\zeta_{j,\ell,1}^{\top}U_{j}\big{]}^{\top}\right\rVert_{\rm F}$ (57)
$\lesssim\kappa_{0}^{2}pr\Delta^{-2},$
with probability at least $1-C_{1}^{\prime}Le^{-c_{1}^{\prime}p}$.
Now we bound $\sum_{\ell=1}^{L}w_{\ell}f_{j,\ell}\big{[}\zeta_{j,\ell,1}U_{j}\big{]}U_{j}^{\top}$. The other three terms can be bounded similarly. Since $\mathrm{vec}(G_{j,\ell}[UV]_{-j,\ell}^{\top}Z_{j,\ell}^{\top}U_{j})=(U_{j}^{\top}\otimes G_{j,\ell}[UV]_{-j,\ell}^{\top})\mathrm{vec}(Z_{j,\ell}^{\top})$, we have that
$\mathrm{vec}(G_{j,\ell}[UV]_{-j,\ell}^{\top}Z_{j,\ell}^{\top}U_{j})\sim\mathcal{N}(0,\sigma_{\ell}^{2}(I_{r_{j,U}}\otimes\Lambda_{j,\ell}^{2})),$
which shows that $G_{j,\ell}[UV]_{-j,\ell}^{\top}Z_{j,\ell}^{\top}U_{j}$ has independent normal entries with variance at most $\lambda_{\max}^{2}\sigma_{\ell}^{2}$. Let $G_{j,\ell}[UV]_{-j,\ell}^{\top}Z_{j,\ell}^{\top}U_{j}=(\bm{w}_{1,j,\ell},\dots,\bm{w}_{r_{j,U},j,\ell})$. Since
$\sum_{\ell=1}^{L}w_{\ell}f_{j,\ell}\big{[}\zeta_{j,\ell,1}U_{j}\big{]}=\sum_{\ell=1}^{L}w_{\ell}f_{j,\ell}([U_{j}V_{j,\ell}]G_{j,\ell}[UV]_{-j,\ell}^{\top}Z_{j,\ell}^{\top}U_{j}),$
we can write
$\sum_{\ell=1}^{L}w_{\ell}f_{j,\ell}\big{[}\zeta_{j,\ell,1}U_{j}\big{]}=\left(\sum_{\ell=1}^{L}w_{\ell}O_{1,j,\ell}[U_{j}V_{j,\ell}]\bm{w}_{1,j,\ell},\dots,\sum_{\ell=1}^{L}w_{\ell}O_{r_{j,U},j,\ell}[U_{j}V_{j,\ell}]\bm{w}_{r_{j,U},j,\ell}\right).$
In the above matrix, each column $\sum_{\ell=1}^{L}w_{\ell}O_{k,j,\ell}[U_{j}V_{j,\ell}]\bm{w}_{k,j,\ell}$ is independent among different $k$ and has norm distribution $\mathcal{N}\left(0,\Sigma_{j,k}\right)$, where
$\Sigma_{j,k}=\sum_{\ell=1}^{L}w_{\ell}^{2}\sigma_{\ell}^{2}O_{1,j,\ell}[U_{j}V_{j,\ell}]\Lambda_{j,\ell}^{2}[U_{j}V_{j,\ell}]^{\top}O_{1,j,\ell}^{\top}.$
Therefore, $\mathrm{vec}\left(\sum_{\ell=1}^{L}w_{\ell}f_{j,\ell}\big{[}\zeta_{j,\ell,1}U_{j}\big{]}\right)\sim\mathcal{N}(0,\Sigma_{j})$, where $\Sigma_{j}=\mathrm{diag}(\Sigma_{j,1},\Sigma_{j,2},\dots,\Sigma_{j,r_{j,U}})$. Since
$\left\lVert\Sigma_{j}\right\rVert_{2}\leq\max_{k}\left\lVert\Sigma_{j,k}\right\rVert_{2}\leq\sum_{\ell=1}^{L}w^{2}_{\ell}\sigma_{\ell}^{2}\lambda^{2}_{\max}\lambda_{\min}^{-2}\Delta^{-2},$
we obtain that
$\left\lVert\sum_{\ell=1}^{L}w_{\ell}f_{j,\ell}\big{[}\zeta_{j,\ell,1}U_{j}\big{]}U_{j}^{\top}\right\rVert_{\rm F}$ $=\left\lVert\sum_{\ell=1}^{L}w_{\ell}f_{j,\ell}\big{[}\zeta_{j,\ell,1}U_{j}\big{]}\right\rVert_{\rm F}$ (58)
$=\left\lVert\mathrm{vec}\left(\sum_{\ell=1}^{L}w_{\ell}f_{j,\ell}\big{[}\zeta_{j,\ell,1}U_{j}\big{]}\right)\right\rVert_{2}$
$\leq\left\lVert\Sigma_{j}^{1/2}\right\rVert_{2}\left\lVert\Sigma_{j}^{-1/2}\mathrm{vec}\left(\sum_{\ell=1}^{L}w_{\ell}f_{j,\ell}\big{[}\zeta_{j,\ell,1}U_{j}\big{]}\right)\right\rVert_{2}$
$\leq\sqrt{2pr}\kappa_{0}\sqrt{\sum_{\ell=1}^{L}w_{\ell}^{2}\sigma_{\ell}^{2}}\Delta^{-1},$
with probability at least $1-2e^{-pr/8}$, where we apply (41) to bound the norm of the vector $\Sigma_{j}^{-1/2}\mathrm{vec}\left(\sum_{\ell=1}^{L}w_{\ell}f_{j,\ell}\big{[}\zeta_{j,\ell,1}U_{j}\big{]}\right)\sim\mathcal{N}(0,I_{p_{j}r_{j,U}})$.
Therefore, by Davis-Kahan Theorem, we obtain
$\quad\sup_{j}\left\lVert\widehat{U}_{j}\widehat{U}_{j}^{\top}-U_{j}U_{j}^{\top}\right\rVert_{\rm F}$
$\leq 2\left\lVert\sum_{\ell=1}^{L}w_{\ell}\widehat{U}_{j,\ell}\widehat{U}_{j,\ell}^{\top}-U_{j}U_{j}^{\top}\right\rVert_{\rm F}$
@@
$\leq\widetilde{C}_{2}\left(\frac{\sqrt{pr}\kappa_{0}\sqrt{\sum_{\ell=1}^{L}w_{\ell}^{2}\sigma_{\ell}^{2}}}{\Delta}+\frac{pr\kappa_{0}^{2}\sum_{\ell=1}^{L}w_{\ell}\sigma_{\ell}^{2}}{\Delta^{2}}\right),$
with probability at least $1-\widetilde{C}_{1}e^{-\widetilde{c}_{1}p}$.
Using the same argument in the proof of Theorem 3.1, for any $\ell$, we have
$\quad\sup_{j}\left\lVert\widehat{V}_{j,\ell}\widehat{V}_{j,\ell}^{\top}-V_{j,\ell}V_{j,\ell}^{\top}\right\rVert_{\rm F}$
$\lesssim\frac{\sqrt{pr_{V}}\kappa_{0}}{\lambda_{\min}}\left(\sigma_{\ell}+\kappa\kappa_{0}\sqrt{r}\sqrt{\sum_{\ell=1}^{L}w_{\ell}^{2}\sigma_{\ell}^{2}}\right)+\frac{\sqrt{r_{V}}pr\kappa^{2}\kappa_{0}^{2}\sum_{\ell=1}^{L}w_{\ell}\sigma_{\ell}^{2}}{\lambda_{\min}^{2}}.$
with probability at least $1-\widetilde{C}_{1}e^{-\widetilde{c}_{1}p}$.
Since $\kappa_{0},\kappa=O(1)$, $\Delta/\max_{\ell}\sigma_{\ell}\gg\sqrt{prL}$ and
$\sum_{\ell=1}^{L}w_{\ell}\sigma_{\ell}^{2}\leq\sqrt{L}\max_{\ell}\sigma_{\ell}\sqrt{\sum_{\ell=1}^{L}w_{\ell}^{2}\sigma_{\ell}^{2}},$
@@
$\frac{pr\kappa_{0}^{2}\sum_{\ell=1}^{L}w_{\ell}\sigma_{\ell}^{2}}{\Delta^{2}}\lesssim\frac{\sqrt{pr}\kappa_{0}\sqrt{\sum_{\ell=1}^{L}w_{\ell}^{2}\sigma_{\ell}^{2}}}{\Delta},$
and thus, with probability at least $1-\widetilde{C}_{1}e^{-\widetilde{c}_{1}p}$,
$\sup_{j}\left\lVert\widehat{U}_{j}\widehat{U}_{j}^{\top}-U_{j}U_{j}^{\top}\right\rVert_{\rm F}\lesssim\frac{\sqrt{pr}\kappa_{0}\sqrt{\sum_{\ell=1}^{L}w_{\ell}^{2}\sigma_{\ell}^{2}}}{\Delta}.$
Similarly, it holds that
$\frac{\sqrt{r_{V}}pr\kappa^{2}\kappa_{0}^{2}\sum_{\ell=1}^{L}w_{\ell}\sigma_{\ell}^{2}}{\lambda_{\min}^{2}}\lesssim\frac{\sqrt{pr_{V}}\kappa_{0}}{\lambda_{\min}}\kappa\kappa_{0}\sqrt{r}\sqrt{\sum_{\ell=1}^{L}w_{\ell}^{2}\sigma_{\ell}^{2}},$
which yields that, with probability at least $1-\widetilde{C}_{1}e^{-\widetilde{c}_{1}p}$,
$\sup_{j}\left\lVert\widehat{V}_{j,\ell}\widehat{V}_{j,\ell}^{\top}-V_{j,\ell}V_{j,\ell}^{\top}\right\rVert_{\rm F}\lesssim\frac{\sqrt{pr_{V}}\kappa_{0}}{\lambda_{\min}}\Big{(}\sigma_{\ell}+\kappa\kappa_{0}\sqrt{r}\sqrt{\sum_{\ell=1}^{L}w_{\ell}^{2}\sigma_{\ell}^{2}}\Big{)}.$
Plugging in $w_{\ell}=\frac{\sigma_{\ell}^{-2}}{\sum_{\ell=1}^{L}\sigma_{\ell}^{-2}}$ leads to
$\sup_{j}\left\lVert\widehat{U}_{j}\widehat{U}_{j}^{\top}-U_{j}U_{j}^{\top}\right\rVert_{\rm F}\lesssim\frac{\sqrt{pr}\kappa_{0}}{\Delta\sqrt{\sum_{\ell=1}^{L}\sigma_{\ell}^{-2}}},$
and
$\sup_{j}\left\lVert\widehat{V}_{j,\ell}\widehat{V}_{j,\ell}^{\top}-V_{j,\ell}V_{j,\ell}^{\top}\right\rVert_{\rm F}\lesssim\frac{\sqrt{pr_{V}}\kappa_{0}}{\lambda_{\min}}\Big{(}\sigma_{\ell}+\frac{\kappa\kappa_{0}\sqrt{r}}{\sqrt{\sum_{\ell=1}^{L}\sigma_{\ell}^{-2}}}\Big{)}.$
Proof of Theorem C.2.
With adaptive weights $\widehat{w}_{\ell}$, we have, by Davis-Khan Theorem,
$\quad\sup_{j}\left\lVert\widehat{U}_{j}\widehat{U}_{j}^{\top}-U_{j}U_{j}^{\top}\right\rVert_{\rm F}$
$\leq 2\sup_{j}\left\lVert\sum_{\ell=1}^{L}\widehat{w}_{\ell}\widehat{U}_{j,\ell}\widehat{U}_{j,\ell}^{\top}-U_{j}U_{j}^{\top}\right\rVert_{\rm F}$
$\leq 2\sup_{j}\sum_{\ell=1}^{L}\left\lvert\widehat{w}_{\ell}-w_{\ell}^{*}\right\rvert\left\lVert\widehat{U}_{j,\ell}\widehat{U}_{j,\ell}^{\top}-U_{j}U_{j}^{\top}\right\rVert_{\rm F}$
$\quad+2\sup_{j}\left\lVert\sum_{\ell=1}^{L}w_{\ell}^{*}\widehat{U}_{j,\ell}\widehat{U}_{j,\ell}^{\top}-U_{j}U_{j}^{\top}\right\rVert_{\rm F}.$
By the proof of Theorem C.1, with probability at least $1-\widetilde{C}_{1}e^{-\widetilde{c}_{1}p}$,
$\sup_{j}\left\lVert\sum_{\ell=1}^{L}w_{\ell}^{*}\widehat{U}_{j,\ell}\widehat{U}_{j,\ell}^{\top}-U_{j}U_{j}^{\top}\right\rVert_{\rm F}\lesssim\frac{\sqrt{pr}\kappa_{0}}{\Delta\sqrt{\sum_{\ell=1}^{L}\sigma_{\ell}^{-2}}}.$
Also, by (49) and Davis-Khan Theorem, with probability at least $1-C_{1}^{\prime}Le^{-c_{1}^{\prime}p}$,
$\sup_{j,\ell}\frac{1}{\sigma_{\ell}}\left\lVert\widehat{U}_{j,\ell}\widehat{U}_{j,\ell}^{\top}-U_{j}U_{j}^{\top}\right\rVert_{\rm F}\lesssim\frac{\sqrt{pr}\kappa_{0}}{\Delta}.$
Now we bound $\left\lvert\widehat{w}_{\ell}-w_{\ell}^{*}\right\rvert$. Recall that
$\widehat{\sigma}_{\ell}=\Big{\|}\mathcal{T}_{\ell}-\mathcal{T}_{\ell}\times_{1}\big{[}\widehat{U}_{1,\ell}\;\widehat{V}_{1,\ell}\big{]}\big{[}\widehat{U}_{1,\ell}\;\widehat{V}_{1,\ell}\big{]}^{\top}\times_{2}\cdots\times_{J}\big{[}\widehat{U}_{J,\ell}\;\widehat{V}_{J,\ell}\big{]}\big{[}\widehat{U}_{J,\ell}\;\widehat{V}_{J,\ell}\big{]}^{\top}\Big{\|}_{\rm F}/\sqrt{p_{1}p_{2}\cdots p_{J}}.$
By the proof of Lemma 1 in Xia et al. (2022), we have that,
$\left\lvert\widehat{\sigma}_{\ell}\sqrt{p_{1}p_{2}\cdots p_{J}}-\left\lVert\mathcal{Z}_{\ell}\right\rVert_{\rm F}\right\rvert\leq C^{\prime}_{2}\sqrt{pr}\kappa_{0}\sigma_{\ell},\quad\forall\ell,$
with probability $1-C_{1}^{\prime}Le^{-c_{1}^{\prime}p}$.
By Lemma 1 of Laurent and Massart (2000), it holds that
$\mathbb{P}\left(\left\lvert\frac{\left\lVert\mathcal{Z}_{\ell}\right\rVert_{\rm F}^{2}}{\sigma_{\ell}^{2}}-p_{1}p_{2}\cdots p_{J}\right\rvert\geq 2\sqrt{p_{1}p_{2}\cdots p_{J}}\sqrt{c_{3}+1}\sqrt{\log p}+2(c_{3}+1)\log p\right)\leq 2p^{-c_{3}-1},$
and thus, since $L=O(p^{c_{3}})$,
$\sup_{\ell}\left\lvert\frac{\left\lVert\mathcal{Z}_{\ell}\right\rVert_{\rm F}^{2}}{\sigma_{\ell}^{2}}-p_{1}p_{2}\cdots p_{J}\right\rvert\lesssim p^{J/2}\sqrt{\log p},$
with probability at least $1-\widetilde{C}_{1}p^{-1}$ for some constant $\widetilde{C}_{1}$. Hence,
$\left\lvert\left\lVert\mathcal{Z}_{\ell}\right\rVert_{\rm F}-\sigma_{\ell}\sqrt{p_{1}p_{2}\cdots p_{J}}\right\rvert\leq\frac{\left\lvert\left\lVert\mathcal{Z}_{\ell}\right\rVert_{\rm F}^{2}-\sigma_{\ell}^{2}p_{1}p_{2}\cdots p_{J}\right\rvert}{\left\lVert\mathcal{Z}_{\ell}\right\rVert_{\rm F}+\sigma_{\ell}\sqrt{p_{1}p_{2}\cdots p_{J}}}\lesssim\sigma_{\ell}\sqrt{\log p},$
@@
$\left\lvert\widehat{\sigma}_{\ell}-\sigma_{\ell}\right\rvert\leq\widetilde{C}_{2}p^{-(J-1)/2}\sqrt{r}\kappa_{0}\sigma_{\ell},$
with probability at least $1-\widetilde{C}_{1}p^{-1}$ for some constants $\widetilde{C}_{1}$ and $\widetilde{C}_{2}$. This further leads to
$\frac{(1-\widetilde{C}_{2}p^{-(J-1)/2}\kappa_{0}\sqrt{r})^{2}}{(1+\widetilde{C}_{2}p^{-(J-1)/2}\kappa_{0}\sqrt{r})^{2}}\frac{\sigma_{\ell}^{-2}}{\sum_{\ell=1}^{L}\sigma_{\ell}^{-2}}\leq\widehat{w}_{\ell}=\frac{\widehat{\sigma}_{\ell}^{-2}}{\sum_{\ell=1}^{L}\widehat{\sigma}_{\ell}^{-2}}\leq\frac{(1+\widetilde{C}_{2}p^{-(J-1)/2}\kappa_{0}\sqrt{r})^{2}}{(1-\widetilde{C}_{2}p^{-(J-1)/2}\kappa_{0}\sqrt{r})^{2}}\frac{\sigma_{\ell}^{-2}}{\sum_{\ell=1}^{L}\sigma_{\ell}^{-2}},$
and thus
$\left\lvert\widehat{w}_{\ell}-w^{*}_{\ell}\right\rvert\leq C_{W}p^{-(J-1)/2}\kappa_{0}\sqrt{r}\frac{\sigma_{\ell}^{-2}}{\sum_{\ell=1}^{L}\sigma_{\ell}^{-2}},\forall\ell,$
for some constant $C_{W}$, with probability at least $1-\widetilde{C}_{1}p^{-1}$. Therefore,
$\sup_{j}\sum_{\ell=1}^{L}\left\lvert\widehat{w}_{\ell}-w_{\ell}^{*}\right\rvert\left\lVert\widehat{U}_{j,\ell}\widehat{U}_{j,\ell}^{\top}-U_{j}U_{j}^{\top}\right\rVert_{\rm F}\lesssim p^{-(J-1)/2}\kappa_{0}\sqrt{r}\frac{\sum_{\ell=1}^{L}\sigma_{\ell}^{-1}}{\sum_{\ell=1}^{L}\sigma_{\ell}^{-2}}\frac{\sqrt{pr}\kappa_{0}}{\Delta}\leq\sqrt{\frac{rL}{p^{J-1}}}\frac{\sqrt{pr}\kappa_{0}}{\Delta\sqrt{\sum_{\ell=1}^{L}\sigma_{\ell}^{-2}}},$
with probability at least $1-\widetilde{C}_{1}p^{-1}$. Then we obtain that, with probability at least $1-\widetilde{C}_{1}p^{-1}$,
$\sup_{j}\left\lVert\widehat{U}_{j}\widehat{U}_{j}^{\top}-U_{j}U_{j}^{\top}\right\rVert_{\rm F}\lesssim\frac{\sqrt{pr}\kappa_{0}}{\Delta\sqrt{\sum_{\ell=1}^{L}\sigma_{\ell}^{-2}}}+\frac{r\kappa_{0}\sqrt{L}}{\Delta\sqrt{p^{J-2}\sum_{\ell=1}^{L}\sigma_{\ell}^{-2}}},$